# Correcting Negative Bias in Large Language Models through Negative Attention Score Alignment

### v1

@@ Abstract
A binary decision task, like yes-no questions or answer verification, reflects a significant real-world scenario such as where users look for confirmation about the correctness of their decisions on specific issues. In this work, we observe that language models exhibit a negative bias in the binary decisions of complex reasoning tasks. Based on our observations and the rationale about attention-based model dynamics, we propose a negative attention score (NAS) to systematically and quantitatively formulate negative bias. Based on NAS, we identify attention heads that attend to negative tokens provided in the instructions as answer candidate of binary decisions, regardless of the question in the prompt, and validate their association with the negative bias. Additionally, we propose the negative attention score alignment (NASA) method, which is a parameter-efficient fine-tuning technique to address the extracted negatively biased attention heads. Experimental results from various domains of reasoning tasks and large model search space demonstrate that NASA significantly reduces the gap between precision and recall caused by negative bias while preserving their generalization abilities. Our codes are available at https://github.com/ysw1021/NASA.
@@
Sangwon Yu1thanks: Equal Contribution    Jongyoon Song [^1] footnotemark: 1    Bongkyu Hwang2    Hoyoung Kang2    Sooah Cho2 Junhwa Choi2    Seongho Joe2    Taehee Lee2   Youngjune L. Gwon2   Sungroh Yoon1,3thanks:   Corresponding author 1Department of Electrical and Computer Engineering, Seoul National University 2Samsung SDS, Korea 3Interdisciplinary Program in Artificial Intelligence, Seoul National University {dbtkddnjs96, coms1580, sryoon}@snu.ac.kr {bongkyu.hwang, hoyoung.kang, sooah.cho, jh496.choi, drizzle.cho, taehee23.lee, gyj.gwon}@samsung.com
@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled complex reasoning tasks executed by understanding user instructions (Ouyang et al., 2022; Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2023). As the capabilities of LLMs have expanded rapidly, research has intensified to analyze their characteristics and inherent issues. One of the major issues is the generation of factually incorrect content, known as “hallucination”, which significantly degrades the reliability of LLM-based services (Zhang et al., 2023; Xu et al., 2024; Huang et al., 2023). The hallucination problem in LLMs can be attributed to factors such as parametric knowledge, overconfidence, and biases (Zhang et al., 2023). Studies have been conducted to understand the decision-making mechanisms in LLMs from a parameter perspective. For instance, the “logit lens” technique allows for the interpretation of the model’s reasoning process along the layers, exploiting the hidden representations from intermediate layers (Belrose et al., 2023; Ferrando et al., 2023). This technique has been utilized to analyze model characteristics in various scenarios such as in-context learning. On the other hand, Yuan et al. (2024) focus on analyzing hallucination phenomenon based on attention heads. However, the causes and mechanisms of hallucination in LLMs vary by type of task, indicating that extensive further study is still necessary.
@@
In this paper, we aim to identify biases that emerge when LLMs respond to questions requiring a binary decision such as affirmation or negation. Specifically, in yes-no question-answering (QA) tasks that demand complex reasoning, such as mathematical or logical reasoning, we observe a negative bias in LLMs. This negative bias, which implies a discrepancy between the model’s reasoning ability and accuracy, potentially decreases the reliability of model predictions. From the rationale that the model allocates higher attention to the negative candidates among the binary answer options presented in the user prompt, we propose a negative attention score (NAS) to systematically probe these negatively biased attention heads in large language models (LLMs). Based on NAS, we detect attention heads that attend to tokens associated with negation in an instruction. Our findings also confirm the existence of attention heads that are predominantly involved in manifesting this negative bias.
To address the identified attention heads, we introduce a parameter-efficient fine-tuning technique named NAS alignment (NASA). In NASA, the model is incrementally fine-tuned, starting with the most biased attention head. The method includes periodic monitoring of NAS which is used for automatic scheduling of early stopping and update cancellation. Experimental results demonstrate that our method significantly reduces negative bias while maintaining general reasoning performance. Additionally, we observe an improvement in calibration for yes-no QA tasks in LLMs.
@@
  • We observe that LLMs exhibit a negative bias in binary decisions requiring complex reasoning and demonstrate its association with attention heads.
  • We propose NAS, a metric and framework that allows for systematic probing of attention heads involved in negative bias.
  • Based on NAS, we introduce NASA, a parameter-efficient fine-tuning technique that successfully addresses negative bias.
@@ 2 Related Work
2.1 Hallucination Problems in Complex Reasoning Tasks
LLMs are demonstrating capabilities not only in typical natural language processing tasks but also in tasks that require complex thought processes. In particular, the use of emergent abilities, such as the chain-of-thoughts, has significantly enhanced the performance of complex reasoning (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Madaan et al., 2023). However, the increased complexity of tasks and pipelines has made the hallucination problem in complex reasoning tasks challenging to analyze and resolve, which remains an active area of research (Zhang et al., 2023). With the advent of benchmarks vulnerable to hallucinations, recent LLMs are laying the groundwork for analyzing the causes of LLM hallucinations based on substantial data. For instance, StrategyQA (Geva et al., 2021) and MuSiQue (Trivedi et al., 2022) involve queries and contexts that demand multi-step reasoning from LLMs, while GSM8k (Cobbe et al., 2021) involves queries that require mathematical reasoning skills.
In this paper, we aim to understand hallucinations from the perspective of model parameters, using the currently employed complex reasoning datasets. Specifically, we focus on binary decision tasks involving affirmation and negation to identify new characteristics of the model.
2.2 Analysis on Roles and Knowledge in Model Parameters
As the capacity and complexity of LLMs increase, the need to understand and interpret these models is becoming more crucial. In particular, there is active research into analyzing parametric knowledge and model behavior that influence the phenomenon of model hallucination. Meng et al. (2022) propose a framework to specify the location where the knowledge of the model is stored, while Belrose et al. (2023) and Yang et al. (2024) research to interpret the inference process of the model. Yuan et al. (2024) suggest a method to detect model parameters involved in the false premise hallucination phenomenon of LLMs and address this issue.
Similarly, we demonstrate that the model’s negative bias is related to the attention heads. Specifically, we explore the existence and characteristics of query-agnostic negative bias in attention heads. Furthermore, we demonstrate that this can be significantly addressed through a parameter-efficient fine-tuning method.
3 Negative Bias in Large Language Models
A binary decision task, such as yes-no QA or answer verification, represents a major real-world scenario where users seek confirmation on whether their decisions regarding specific problems are correct. This acts as the basic block of the high-dimensional reasoning tasks that need step-by-step sub-tasks. Despite the importance of the binary decision task, unfortunately, many of the complex reasoning benchmark sets are not in the form of binary decisions. Therefore, in this section, we first transform existing complex reasoning tasks into yes-no binary decision tasks. We then conduct statistical observations and quantitative analyses on how various large language models perform in these binary decision scenarios.
@@ 3 Negative Bias in Large Language Models
3.1 Statistical Observation on Complex Reasoning Tasks
@@ 3.1.1 Setup for Binary Decision Task
In this work, we employ two techniques to transform general QA reasoning tasks into binary decision tasks. We modify each sample into a positive or negative sample where the answer is “Yes” or “No”, respectively. In this section, we focus on the case of short-answer QA datasets. We refer to Appendix C for the case of multiple choice datasets.
For the positive sample, we perform a rule-based transformation on general queries by appending a confirmatory query. For example, the query "What is 1+1?" is transformed into "What is 1+1? Is the answer 2?" This approach is straightforward but is limited to cases where the correct answer is “Yes”. Formally, given the question and its label, we transform the prompt for the binary decision task whose target answer is “Yes”:
Prompt I (Binary Decision Transformation) You are given a question and you MUST answer Yes or No. Question: $\{question\}$ Is the answer $\{label\}$? Answer:
To transform into the negative sample, a wrong answer is necessary based on the query and answer which will be placed in label in Prompt I. For this, we obtain suitable incorrect answers by prompting GPT-4. Subsequently, we use the aforementioned rule-based method to convert the original QA samples into binary decision samples where the answer is “No”. Details of GPT-4 prompts can be found in Appendix C.
Figure 3: An example of the negative attention head. Three queries in the figure are sampled from StrategyQA, GSM8K, and MATH datasets.
@@
We use StrategyQA, MuSiQue, subsets of MetaMATH (GSM8k-Rephrased and MATH-Rephrased; Yu et al., 2023), and AR-LSAT for evaluation, covering three domain tasks: multi-hop open-domain QA, mathematical reasoning, and logical reasoning. StrategyQA is a yes-no QA dataset, MuSiQue and the rephrased versions of GSM8k and MATH are short-answer QA datasets, and AR-LSAT is a multiple-choice dataset. Examples of each dataset can be found in Table 12.
To assess reasoning capabilities in binary decision tasks, we convert these datasets into binary decision sets using the aforementioned techniques, except for StrategyQA. Specifically, AR-LSAT is converted to the straightforward yes-no QA formats, while the others are rephrased for answer verification. Details on the conversion to binary decision datasets can be found in Appendix C.
We sample a portion of the entire dataset, and detailed statistics can be found in Appendix A. For the search space of LLMs, we explore the dynamics of LLaMA3-8B-Instruct, Mistral-7B-Instruct-v0.3, Gemma-1.1-7b-it, Qwen2-7B-Instruct, and GPT-4.
@@ 3.1.2 Observation of Negative Bias
Fig 1 presents the precision and recall results for models across various datasets. In most cases, a significant drop in recall compared to precision is observed. As defined, high precision indicates that a high frequency of the model’s positive responses (TP+FP) are positive cases (TP). High recall means that among all the positive cases (TP+FN), the model frequently responds correctly with positive answers (TP). Therefore, intuitively, the phenomenon where existing models exhibit high precision but low recall can be explained to that the model is overly cautious in outputting positive responses. Conversely, this implies that the model outputs negative responses indiscriminately, indicating that the trustworthiness of negative responses and positive responses is not balanced.
Fig 2 shows a histogram of the response confidence of the LLaMA3 model in mathematical and logical reasoning tasks. The X-axis denotes confidence for all lines. Each line represents FP (false positive), FN (false negative), TP (true positive), and TN (true negative). We observe that the overall frequency and confidence of negative responses, including FN and TN, are higher than those of positive responses, including FP and TP. This indicates that the model tends to output negative responses more frequently and with greater confidence.
In summary, through statistical observation, we find that LLMs typically show 1) over-cautiousness to the positive response, and 2) more frequent and more confident negative response. Thus, it is evident that large language models possess a bias towards negative responses in the binary decision-making process of complex reasoning tasks. Detailed case study of the negative responses about the negative bias can be found in Appendix E.
@@ 3.2.1 NAS: Negative Attention Score
In this section, we formulate the previously discovered negative bias from the perspective of model intrinsic properties. Specifically, we focus on the internal attention patterns of the model. Generally, to obtain an answer for a binary decision, a user provides answer candidates like “Yes” or “No” in the form of instructions to the model before posing the query prompt. The model follows these instructions and responds to the given query in the manner provided in the instructions. During this response process, the model attends to the given instructions, and due to the operational characteristics of attention-based models, the candidate with the larger attention weight generally appears in the response. In this context, the negative bias of the LLM can be seen as originating from assigning greater attention weight to negative answer candidates during the reasoning process. Fig 3 illustrates an example of such a "negative head". To validate this rationale, we define the following Negative Attention Score (NAS) using the attention weights applied to negative answer candidate tokens like “No”, and positive answer candidate tokens like “Yes”, observed in a attention head.
Formally, let $L_{P}$ be the length of the prompt $x$, $t_{Yes}$ and $t_{No}$ the positions of the “Yes” and “No” tokens within the instruction, and $L_{I}$ the length of the instruction. For the attention weight inferred by the $h^{th}$ attention head in the $l^{th}$ layer, denoted as $A^{l,h}\in\mathbb{R}^{L_{P}\times L_{P}}$, the NAS is defined as:
@@
$\sum^{L_{P}}_{i=L_{I}}{\left(A^{l,h}_{i,t_{Yes}}+A^{l,h}_{i,t_{No}}\right)*\log\left(\frac{A^{l,h}_{i,t_{No}}}{A^{l,h}_{i,t_{Yes}}}\right)},$ (1)
where we omit the input $x$ in the attention function for brevity. In the $i$-th row, each NAS term consists of two factors: the sum of the attention weights applied to the “Yes” and “No” tokens, which aims to find attention heads focusing significantly on the tokens representing answer options; and the logarithm of the ratio of attention weights applied to these tokens, identifying heads that preferentially attend to the “No” token over the “Yes” token.
Basically, NAS can have a value for each attention head on a single sample. Based on this, we introduce two types of NAS variants: single head NAS and model NAS. Single head NAS represents the average NAS across a specific attention head for the sample or dataset. It is used to quantify the negative bias of a single attention. On the other hand, model NAS represents the total NAS across all attention heads for the sample or dataset. This is used to approximate the model’s own negative bias. For example, we can formally express the single head NAS of $h$-th attention head in $l$-th layer on the dataset $X$ as $\frac{1}{|X|}\sum_{x_{i}\in X}\text{NAS}^{l,h}_{x_{i}}$.

### v2

@@ Abstract
A binary decision task, like yes-no questions or answer verification, reflects a significant real-world scenario such as where users look for confirmation about the correctness of their decisions on specific issues. In this work, we observe that language models exhibit a negative bias in the binary decisions of complex reasoning tasks. Based on our observations and the rationale about attention-based model dynamics, we propose a negative attention score (NAS) to systematically and quantitatively formulate negative bias. Based on NAS, we identify attention heads that attend to negative tokens provided in the instructions as answer candidate of binary decisions, regardless of the question in the prompt, and validate their association with the negative bias. Additionally, we propose the negative attention score alignment (NASA) method, which is a parameter-efficient fine-tuning technique to address the extracted negatively biased attention heads. Experimental results from various domains of reasoning tasks and large model search space demonstrate that NASA significantly reduces the gap between precision and recall caused by negative bias while preserving their generalization abilities.
@@
Sangwon Yu1thanks: Equal Contribution    Jongyoon Song [^1] footnotemark: 1    Bongkyu Hwang2    Hoyoung Kang2    Sooah Cho2 Junhwa Choi2    Seongho Joe2    Taehee Lee2   Youngjune Gwon2   Sungroh Yoon1,3thanks:   Corresponding author 1Department of Electrical and Computer Engineering, Seoul National University 2Samsung SDS, Korea 3AIIS, ASRI, INMC, ISRC, and IPAI, Seoul National University {dbtkddnjs96, coms1580, sryoon}@snu.ac.kr {bongkyu.hwang, hoyoung.kang, sooah.cho, jh496.choi, drizzle.cho, taehee23.lee, gyj.gwon}@samsung.com
@@ 1 Introduction
Recent advancements in large language models (LLMs) have enabled complex reasoning tasks executed by understanding user instructions (Ouyang et al., 2022; Touvron et al., 2023; Achiam et al., 2023; Jiang et al., 2023). As the capabilities of LLMs have expanded rapidly, research has intensified to analyze their characteristics and inherent issues. One of the major issues is the generation of factually incorrect content, known as “hallucination”, which significantly degrades the reliability of LLM-based services (Zhang et al., 2023; Xu et al., 2024; Huang et al., 2023).
The hallucination problem in LLMs can be attributed to factors such as parametric knowledge, overconfidence, and biases (Zhang et al., 2023). Studies have been conducted to understand the decision-making mechanisms in LLMs from a parameter perspective. For instance, the “logit lens” technique allows for the interpretation of the model’s reasoning process along the layers, exploiting the hidden representations from intermediate layers (Belrose et al., 2023; Ferrando et al., 2023). This technique has been utilized to analyze model characteristics in various scenarios such as in-context learning. On the other hand, Yuan et al. (2024) focus on analyzing hallucination phenomenon based on attention heads. However, the causes and mechanisms of hallucination in LLMs vary by type of task, indicating that extensive further study is still necessary.
In this paper, we aim to identify biases that emerge when LLMs respond to questions requiring a binary decision, such as affirmation or negation. We observe a negative bias in LLMs in yes-no question-answering (QA) and answer verification tasks that demand complex reasoning, such as mathematical or logical reasoning. Specifically, we observe that LLMs generally exhibit high precision but low recall in binary decision tasks. Additionally, we find that the prediction confidence for negative decisions is significantly higher than for positive decisions. This suggests that LLMs tend to be overly cautious when making positive decisions, leading to a phenomenon where the model excessively favors negative decisions as a form of shortcut. This negative bias, which indicates a discrepancy between the model’s reasoning ability and accuracy, may ultimately reduce the reliability of its predictions.
@@
From the rationale that the model allocates higher attention to the negative candidates among the binary answer options presented in the user prompt, we propose a negative attention score (NAS) to systematically probe these negatively biased attention heads in LLMs. Based on NAS, we detect attention heads that attend to tokens associated with negation in an instruction. Our findings also confirm the existence of attention heads that predominantly contribute to manifesting this negative bias, regardless of the given query. To address the identified attention heads, we introduce a parameter-efficient fine-tuning technique named NAS alignment (NASA).
In the NASA framework, we first construct a probing set in the form of a binary decision task to extract negative attention heads. This probing set is derived from an existing short-answer QA dataset. Next, we apply an incremental fine-tuning process to the negative attention heads identified based on NAS. During this process, each head undergoes fine-tuning using the probing set, while NASA includes periodic monitoring of NAS, which automatically schedules early stopping and update cancellations. This pipeline ensures that tuning is conducted only up to an appropriate extent and in the correct sequence, effectively mitigating negative bias without compromising the model’s existing capabilities.
For the evaluation of NASA, we measure the performance of four LLMs: LLaMA3-8B-Instruct (Touvron et al., 2023), Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Gemma-1.1-7b-it (Team et al., 2024), and Qwen2-7B-Instruct (Yang et al., 2024a) as the search space. We conduct a comprehensive evaluation across multiple reasoning tasks, including multi-hop QA: StrategyQA (Geva et al., 2021), MuSiQue (Trivedi et al., 2022), mathematical reasoning: GSM8k-Rephrased and MATH-Rephrased subset in MetaMATH (Yu et al., 2023), and logical reasoning: AR-LSAT (Zhong et al., 2021).
Experimental results demonstrate that our method NASA significantly reduces the gap between precision and recall while maintaining or even improving accuracy and F1 score. This indicates that the NASA method successfully mitigates negative bias without compromising overall performance. Additionally, we confirm that the NASA-tuned model effectively preserves general reasoning performance beyond binary decision-making and exhibits improvements in calibration. Furthermore, NASA demonstrates robust performance across various prompting formats, including few-shot settings.
@@
  • We observe that LLMs exhibit a negative bias in binary decision tasks requiring complex reasoning and demonstrate its association with attention heads.
  • We propose negative attention score (NAS), a metric and framework that allows for systematic probing of attention heads involved in negative bias regardless of the input query.
  • Based on NAS, we introduce NAS alignment (NASA), a parameter-efficient fine-tuning technique that effectively addresses negative bias by tuning the extracted negative attention heads in a proper order.
@@ 2 Related Work
Hallucination problems in complex reasoning tasks The emergent abilities of LLMs, such as the chain-of-thoughts, have significantly enhanced the performance of complex reasoning (Wei et al., 2022; Kojima et al., 2022; Wang et al., 2022; Madaan et al., 2023). However, the increased complexity of tasks and pipelines has made the hallucination problem in complex reasoning tasks challenging to analyze and resolve, which remains an active area of research (Zhang et al., 2023). With the advent of benchmarks vulnerable to hallucinations, recent LLMs are laying the groundwork for analyzing the causes of LLM hallucinations based on substantial data which requires complex reasoning abilities (Geva et al., 2021; Trivedi et al., 2022;Cobbe et al., 2021).
Analysis on roles and knowledge in model parameters As the capacity and complexity of LLMs increase, the need to understand and interpret these models is becoming more crucial. In particular, there is active research into analyzing parametric knowledge and model behavior that influence the phenomenon of model hallucination. Meng et al. (2022) propose a framework to specify the location where the knowledge of the model is stored, while Belrose et al. (2023) and Yang et al. (2024b) research to interpret the inference process of the model. Yuan et al. (2024) suggest a method to detect model parameters involved in the false premise hallucination phenomenon of LLMs.
@@
Recent studies have shown that attention heads in language models play a variety of roles such as knowledge recalling (Zheng et al., 2024b). For example, Jin et al. (2024) find the existence of attention heads in LLMs that either recall parametric knowledge or retrieve information from external contexts and García-Carrasco et al. (2024) propose a method to identify attention heads that increase vulnerability to adversarial attacks.
In this paper, we identify the negative bias in LLMs that leads to hallucinations in binary decision tasks requiring complex reasoning. We focus on the attention heads in LLMs and propose a framework to detect and address the attention heads responsible for the negative bias.
3 Negative Bias in LLMs
A binary decision task, such as yes-no QA or answer verification, represents a major real-world scenario where users seek confirmation on whether their decisions regarding specific problems are correct. For instance, a user may inquire whether a statement formed through various contexts is true. Also, a user may seek validation of the correctness of their proposed solutions to problems requiring complex logical reasoning. Such fundamental interactions, represented as yes-no decisions, act as the basic block of the high-dimensional reasoning tasks that need step-by-step sub-tasks. Despite this significance of the binary decision task, many complex reasoning benchmark datasets are not structured in a binary decision format. In this section, we first transform existing complex reasoning tasks into yes-no binary decision tasks. We then conduct statistical observations and quantitative analyses to examine how LLMs perform in these binary decision scenarios, observing a meaningful negative bias in their decisions. Following this, we propose a systematic framework, the Negative Attention Score (NAS), to formulate and quantify the negative bias exhibited by LLMs.
3.1 Statistical Observation
@@ 3.1.1 Setup for Binary Decision Task
We transform general QA reasoning tasks into binary decision tasks by modifying each sample into a positive or negative sample where the answer is “Yes” or “No”, respectively. In this section, we focus on the case of short-answer QA datasets. We refer to Appendix D for the multiple choice datasets.
For the positive sample, we perform a rule-based transformation on general queries by appending a confirmatory query. For example, the query "What is 1+1?" is transformed into "What is 1+1? Is the answer 2?". This approach is straightforward for the cases where the correct answer is “Yes”. Formally, given the question and its label, we use Prompt I to transform the prompt into the positive sample:
Prompt I (Positive Sample Transformation) You are given a question and you MUST answer Yes or No. Question: $\{question\}$ Is the answer $\{label\}$? Answer:
To transform into the negative sample, a wrong label is necessary based on the query and answer which will be placed in label in Prompt I. For this, we obtain suitable incorrect answers by prompting GPT-4 (OpenAI, 2023). We present GPT-4 prompts for the wrong label generation in Appendix D.
Figure 3: An example of the negative attention head. Three queries in the figure are sampled from StrategyQA (Geva et al., 2021), rephrased GSM8K and MATH datasets (Yu et al., 2023).
@@
We utilize StrategyQA, MuSiQue, subsets of MetaMATH (GSM8k-Rephrased and MATH-Rephrased), and AR-LSAT for evaluation, covering three domain tasks: multi-hop open-domain QA, mathematical reasoning, and logical reasoning. StrategyQA is a yes-no QA dataset, MuSiQue and the rephrased versions of GSM8k and MATH are short-answer QA datasets, and AR-LSAT is a multiple-choice dataset. Samples are in Table 14. To assess reasoning capabilities in binary decision tasks, we convert these datasets into binary decision sets using the techniques above, except for StrategyQA. AR-LSAT is converted to the yes-no QA formats, while the others are rephrased for answer verification formats. Details on the conversion can be found in Appendix D. We sample a portion of the entire dataset, and detailed statistics can be found in Appendix A. For the search space of LLMs, we explore LLaMA3-8B-Instruct (LLaMA), Mistral-7B-Instruct-v0.3 (Mistral), Gemma-1.1-7b-it (Gemma), Qwen2-7B-Instruct (Qwen), and GPT-4.
@@ 3.1.2 Observation of Negative Bias
Fig 1 presents the precision and recall results on the transformed datasets. We use FP, FN, TP, and TN to denote false positive, false negative, true positive, and true negative, respectively. In most cases, a significant drop in recall compared to precision is observed. According to the definition, a high precision implies that a significant proportion of the samples predicted as positive by the model (TP+FP) are true positives. Low recall means that among all the positive samples (TP+FN), the model rarely responds correctly with positive answers (TP). The phenomenon where existing models exhibit high precision but low recall can be explained to that the model is overly cautious in outputting positive responses. Conversely, this implies that the model outputs negative responses indiscriminately, indicating that the trustworthiness of negative responses and positive responses is not balanced.
Fig 2 shows a histogram of the response confidence of the LLaMA in mathematical and logical reasoning tasks. We observe that the overall frequency and confidence of negative responses (FN and TN) tend to be higher than those of positive responses (FP and TP). This indicates that the model tends to output negative responses more frequently and with greater confidence.
In summary, we find that LLMs typically show 1) over-cautiousness to the positive response, and 2) more frequent and confident negative responses. It is evident that large language models possess a bias towards negative responses in the binary decision-making process of complex reasoning tasks. A case study of the negative responses about the negative bias can be found in Appendix G.
@@ 3.2.1 NAS: Negative Attention Score
In this section, we formulate negative bias from the perspective of model intrinsic properties. Specifically, we focus on the model’s internal attention patterns. To obtain an answer for a binary decision, a user generally provides answer candidates like “Yes” or “No” as instructions to the model before posing the query prompt (c.f., Prompt I). The model follows these instructions and responds to the given query in the manner provided in the instructions. During this response process, the model attends to the given instructions, and due to the operational characteristics of attention-based models, the candidate with the larger attention weight generally appears in the response. In this context, the negative bias of the LLM can be seen as originating from assigning greater attention weight to negative answer candidates during the reasoning process. Fig 3 illustrates an example of negative attention head. To validate this rationale, we define Negative Attention Score (NAS) using the attention weights applied to negative answer candidate tokens like “No”, and positive answer candidate tokens like “Yes”, observed in a attention head.
Let $L_{x}$ be the length of the sample $x$, $t_{Yes}$ and $t_{No}$ the positions of the “Yes” and “No” tokens within the instruction, and $L_{I}$ the length of the instruction in $x$. For the attention weight inferred by the $h$-th attention head in the $l$-th layer, denoted as $A^{l,h}\in\mathbb{R}^{L_{x}\times L_{x}}$, the NAS is defined as:
@@
$\sum^{L_{x}}_{i=L_{I}}{\left(A^{l,h}_{i,t_{Yes}}+A^{l,h}_{i,t_{No}}\right)*\log\left(\frac{A^{l,h}_{i,t_{No}}}{A^{l,h}_{i,t_{Yes}}}\right)},$ (1)
where we omit the input $x$ in the attention function for brevity. The NAS term consists of two factors: the sum of the attention weights applied to the “Yes” and “No” tokens, which aims to find attention heads focusing significantly on the tokens representing answer options; and the logarithm of the ratio of attention weights applied to these tokens, identifying heads that preferentially attend to the “No” token over the “Yes” token.
NAS can have a value for each attention head on a single sample. Based on this, we introduce two types of NAS variants: single head NAS and model NAS. Single head NAS represents the average NAS of a specific attention head across the sample(s). It is used to quantify the negative bias of a single attention head. On the other hand, model NAS represents the total NAS across all attention heads for the sample(s). This is used to approximate the model’s negative bias. Formally, given a sample set $X$, we define the single head NAS of the $h$-th attention head in the $l$-th layer and model NAS as follows:
[…]
where $L$ and $H$ are the sets of all layer and attention head indices, respectively.